1. Does the equation 2x + 7 – 3x = –x + 12 has infinitely many solutions, one solution, or no
solutions? Show your work to illustrate your answer.

Answer by robertb(5830)

2x + 7 – 3x = –x + 12
<==> -x + 7 = -x + 12
<==> 7 = 12, a contradiction.
Hence the solution set is the empty set.